Two airplanes leave an airport at the same
time. The velocity of the first airplane is
730 m/h at a heading of 44.3◦. The velocity
of the second is 590 m/h at a heading of 83◦.
How far apart are they after 2.9 h?
Answer in units of m.
1
Expert's answer
2019-10-08T10:07:48-0400
"d_1=730(2.9)=2117\ km"

"d_2=590(2.9)=1711\ km"

"d=\sqrt{2117^2+1711^2+2(2117)(1711)\cos{(83-44.3)\degree}}=3614\ m"